OpenStudy (anonymous):

2b^3-11b+5 please explain step by step if i have the time im having trouble with these

OpenStudy (anonymous):

factoring

OpenStudy (anonymous):

I don't think that can be factored

OpenStudy (anonymous):

Do you mean 2b^2 for the first term

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

(2x - 1)(x-5)

OpenStudy (anonymous):

If it is 2b^2, then \[2b^2-11b+5=2b^2-10b-b+5=2b(b-5)-(b-5)=(b-5)(2b-1)\]

OpenStudy (anonymous):

wow i still dont understand lol thanks anyway